cot(-8*pi/3)
it needs to be done without a calculator

Answer by stanbon(75887)

cot(-8*pi/3)
--
= cot(-2(pi)-(2/3)pi)
---
= cot(-(2/3)pi)
---
= cot (pi/3)
----
= sin(pi/3)/cos(pi/3)
----
= [sqrt(3)/2)]/[1/2)
---
= sqrt(3)
=================
Cheers,
Stan H.